Solve: v/2 – v/3 = 2

: Solve: v/2 – v/3 = 2


Answer by mangopeeler07(442)

v/2 – v/3 = 2

Common denominator=6
3v/6-2v/6=2

Subtract the numerators
v/6=2

Multiply both sides by 6
v=12